The perimeter of a rectangle is 136 inches. If the length is four less than five times the width, what is the area of the rectangle? Thanks for your help.
Answer by macston(5194)

.
W=width; L=length=5W-4; P=perimeter=136 inches
.
P=2(L+W)Substitute for L
136 in=2((5W-4)+W) Divide each side by 2.
68 in=6W-4 Add 4 in to each side.
72 in=6W
12 in=W
.
L=5W-4in=5(12)-4=60-4=56 in
.
Area=LW=(56in)(12in)=672in^2
.
ANSWER: The area is 672 square inches.